• Three lamps are rated 100 W, 50 W, and 10 W If each lamp is connected to a 220-V source, calculate the (a) current through each lamp and (b) resistance through each lamp.
  • The resistance of a bulb is 50Ω. What is the current produced by a 24-V battery in the bulb? How many electrons does this current represent?
1
Expert's answer
2022-04-21T13:01:03-0400

1.

"I=\frac{P}{V},\quad R=\frac{V^2}{P}"

"I_1=\frac{100}{220}=0.45\:{\rm A},\quad R_1=\frac{220^2}{100}=848\:\Omega"

"I_2=\frac{50}{220}=0.23\:{\rm A},\quad R_2=\frac{220^2}{50}=968\:\Omega"

"I_3=\frac{10}{220}=0.05\:{\rm A},\quad R_3=\frac{220^2}{10}=4840\:\Omega"

2.

"I=\frac{V}{R}=\frac{24}{50}=0.48 \:\rm A"

"q=It=0.48\: \rm A*1\:\rm s=0.48 \; C"

"N=\frac{q}{e}=\frac{0.48}{1.6*10^{-19}}=3*10^{18}"